Circular No. 7553

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 Y3 (SCOTTI) Additional astrometry, including prediscovery observations by LIN-

EAR on Nov. 29 and Dec. 21 identified by B. G. Marsden, appear on MPEC 2000-Y47, together with the following orbital elements showing this to be a short-period comet. The elements indicate an approach to within 0.05 AU of Jupiter in Sept. 1998. Further to IAUC 7552, J. V. Scotti notes that the comet showed a 7′′ coma and a 1′.16 tail in p.a. 270^{o} on a Spacewatch CCD image taken on Dec. 31.174 UT. An image obtained at Kleť on Dec. 30.79 shows a coma diameter of 8′′ and m_{1} = 17.5.

T = 2000 Nov. 6.5067 TT  ω = 89.^{o}3025
e = 0.200779  Ω = 355.0977 }2000.0
q = 4.044148 AU  i = 2.2469
a = 5.060113 AU n^{o} = 0.0865892 P = 11.383 years
2000 /01 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 22 6^{h}01.^{m}96 +26^{o}23.^{′}6 3.065 4.048 177.^{o}0 0.^{o}7 17.5
Jan. 1 5 55.30 +26 22.0 3.083 4.050 168.0  2.9 17.5
11 5 49.28 +26 17.8 3.130 4.052 156.6  5.5 17.6
21 5 44.44 +26 11.8 3.206 4.054 145.4  7.9 17.6
31 5 41.18 +26 04.9 3.305 4.057 134.5 10.0 17.7
Feb. 10 5 39.73 +25 58.0 3.424 4.061 124.1 11.6 17.8
20 5 40.15 +25 51.6 3.559 4.064 114.1 12.8 17.8
Mar. 2 5 42.41 +25 46.0 3.704 4.068 104.6 13.6 17.9
12 5 46.38 +25 41.1 3.856 4.072  95.5 14.1 18.0

COMET P/2000 S4 (LINEAR-SPACEWATCH) The IAU Committee on Small Bodies Nomenclature has given to comet

P/2000 S4 (cf. IAUC 7502) the name LINEAR-Spacewatch.

POSSIBLE NOVA IN PUPPIS T. Kato, Kyoto University, forwards the following additional prediscov-

ery photographic magnitudes for the new variable reported on IAUC 7552, from T-Max 400 film exposures obtained by K. Haseda, Aichi, Japan: Nov. 23.78 UT, 8.8; Dec. 1.77, 8.8.

2000 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green